OpenStudy (anonymous):

A school club sold 25 boxes of cookies. Two kinds were sold. The chocalate cookies cost $3.00 a box. The vanilla cookies cost $4.00 per box. The club collected $88.00. How many boxes of each type were sold?

OpenStudy (anonymous):

sorry i just have to get this done by 4:00 and i dont understand it

OpenStudy (anonymous):

i need how u got it too if thats ok

OpenStudy (ash2326):

Let the no. of chocolate cookies boxes sold be x and no. of vanilla cookies boxes sold be y, we 're given total no. of boxes sold is 25 so we get x+y=25

OpenStudy (ash2326):

The price of chocolate cookie box is $3 and vanilla cookie box cost $4 it's given the club collected $88 so 3x+4y=88 we have two equations x+y=25 and 3x+4y=88

OpenStudy (ash2326):

Can you solve the equations to find x and y?

OpenStudy (anonymous):

Ash 2326 do you mind helping me with my problem once you are done?

OpenStudy (ash2326):

Not at all, I'll help you

OpenStudy (anonymous):

i dont really kno how to solve them

OpenStudy (ash2326):

we have 1st.......x+y=25 2nd......3x+4y=88 Multiply Ist by 3 , we get 3x+3y=75 now subtract this from 2nd we get \[(3x+4y)-(3x+3y)=88-75\] \[4y-3y=13\] we get y=13 nox x+y=25 so x+13=25 x=12 12 chocloate cookie boxes 13 vanilla cookie boxes

OpenStudy (anonymous):

so thats the answer?

OpenStudy (ash2326):

yeah, :)

OpenStudy (anonymous):

thank u so much!!

OpenStudy (ash2326):

welcome :)